jackfruits are twice as expensie as watermelons and water melon is one third as
expensive as durians. what ratio of the price of one jackfruit to one durian.

Answer by lwsshak3(6490)

jackfruits are twice as expensie as watermelons and water melon is one third as
expensive as durians. what ratio of the price of one jackfruit to one durian.
**
let x=price of 1 durian
x/3=price of 1 watermelon
2x/3=price of 1 jackfruit
ratio of the price of one jackfruit to one durian=(2x/3)/x=2/3